Company ABC is trying to determine the current value its equity. As of its last financial statements, it had net income of $10,850, with after-tax cash earnings of $1,850. The book value of its equity was $20,000 with cash value of $4,000. The CAPEX was $5,000, depreciation was $2,000, change in working capital was $1,500, and the cash flow from net debt was $1,750. It is assumed at a risk-free rate of 1.5%, a beta of 1.3 and a risk premium of 6%. After a 5-year high growth period, the stable growth will be 2%, but the cost of equity will remain the same. Given this information, what is the equity value of the operating cash flows? Show the calculation in details.
Answer #1

First we calculate cost of equity and free cash flow to equity (FCFE).

Cost of equity = Risk-free rate + beta*market risk premium = 1.5% + 1.3*6% = 1.5% + 7.8% = 9.3%

FCFE = Net income + Depreciation - CAPEX - change in working capital + net borrowing

FCFE = $10,850 + $2,000 - $5,000 - $1,500 + $1,750 = $8,100

FCFE in year 6 = current FCFE*(1+growth rate) = $8,100*(1+0.02) = $8,100*1.02 = $8,262

FCFE value at the end of year 5 = FCFE year 6/(cost of equity - growth rate) = $8,262/(0.093 - 0.02) = $8,262/0.073 = $113,178.08

Equity value today = FCFE value at the end of year 5/(1+growth rate)no. of years = $113,178.08/1.0935 = $113,178.08/1.5599 = $72,554.70